OpenStudy (anonymous):

Question 4 (2 points) Solve 1/2(p+5)-4>1/3(2p-1) Answer: Question 5 (2 points) On the same graph, plot the two functions f(p) = 1/2(p+5)-4 and g(p) = 1/3(2p-1) Include appropriate labels. Explain how this graph relates to the solution of Question 4. Answer:

OpenStudy (anonymous):

I solved Question 4 and I gotLCD as 6 so 3p+15-24>4p-2 3p-9>4p-2 3p-4p>-2+9 -p>7 p<-7

OpenStudy (anonymous):

I solved Question 5 and I got f(p) = (1/2p) + (5/2) - (8/2) = (1/2p) - (3/2) g(p) = (2/3p) -(1/3)

OpenStudy (anonymous):

I can graph it but my question is how do we answer how this graph relates to the solution of Question 4

OpenStudy (anonymous):

Please Help

OpenStudy (anonymous):

Please Help

OpenStudy (anonymous):

Would it be that in Question 4 the solution is in the shaded area and in Question 5 the solution is the shaded area so both of them have similar solutions